### 06 Apr 2026 v1

@@ 3.2 Properties of the reduced SBAL function
$S_{\eta_{\rho}}(\hat{x},s;\mu)[h,h]=\rho\langle h_{\hat{x}},\mathcal{B}^{*}\mathcal{H}^{-1}\mathcal{W}\mathcal{B}h_{\hat{x}}\rangle+\rho^{-1}\langle h_{s},\mathcal{H}^{-1}h_{s}\rangle.$ (52)
Since $\mathcal{B}^{*}$ is surjective and $\mathcal{W}\succ 0$, $S_{\eta_{\rho}}(\hat{x},s;\mu)$ is positive definite. It suffices to prove that
$|D^{3}\eta_{\rho}(\hat{x},s;\mu)[h,h,h]|\leq\dfrac{2}{\sqrt{\mu}}(S_{\eta_{\rho}}(\hat{x},s;\mu)[h,h])^{3/2}.$ (53)
Let $\hat{h}=(\hat{h}_{\hat{x}},\hat{h}_{s})=(\mathcal{H}^{-1}\mathcal{B}h_{\hat{x}},\mathcal{H}^{-1}h_{s})\in\mathbb{E}\times\mathbb{E}$. By Corollary 1 and the formula $\mathcal{H}=\mathcal{I}_{\mathbb{E}}+\mathcal{W}=\mathcal{I}_{\mathbb{E}}+\frac{\mu}{\rho}D^{2}\phi(z_{\rho})$, we have
$S_{\eta_{\rho}}(\hat{x},s;\mu)[h,h]$ (54)
$=$ $\ \rho\langle\hat{h}_{\hat{x}},\mathcal{H}\mathcal{W}\hat{h}_{\hat{x}}\rangle+\rho^{-1}\langle\hat{h}_{s},\mathcal{H}\hat{h}_{s}\rangle$
$=$ $\ \mu D^{2}\phi(z_{\rho})[\hat{h}_{\hat{x}},\hat{h}_{\hat{x}}]+\frac{\mu^{2}}{\rho}\|D^{2}\phi(z_{\rho})\hat{h}_{\hat{x}}\|^{2}+\rho^{-1}\|\hat{h}_{s}\|^{2}+\frac{\mu}{\rho^{2}}D^{2}\phi(z_{\rho})[\hat{h}_{s},\hat{h}_{s}]$
$=$ $\ \mu D^{2}\phi(z_{\rho})[\hat{h}_{\hat{x}}-\rho^{-1}\hat{h}_{s},\hat{h}_{\hat{x}}-\rho^{-1}\hat{h}_{s}]+\dfrac{1}{\rho}\|\mu D^{2}\phi(z_{\rho})\hat{h}_{\hat{x}}+\hat{h}_{s}\|^{2}$
$\geq$ $\ \mu D^{2}\phi(z_{\rho})[\hat{h}_{\hat{x}}-\rho^{-1}\hat{h}_{s},\hat{h}_{\hat{x}}-\rho^{-1}\hat{h}_{s}],$
@@
$D^{2}\eta_{\rho}(\hat{x},s;\mu)[h,h]$ $=\rho\langle\mathcal{B}h_{\hat{x}},(\mathcal{I}_{\mathbb{E}}-\mathcal{H}^{-1})\mathcal{B}h_{\hat{x}}\rangle-\rho^{-1}\langle h_{s},\mathcal{H}^{-1}h_{s}\rangle$ (55)
@@
Direct computation gives
$D_{\hat{x}}\mathcal{H}[h_{\hat{x}},\cdot,\cdot]=\frac{\mu}{\rho}D^{3}\phi(z_{\rho})[\hat{h}_{\hat{x}}$$\cdot,\cdot$], (56a)
$D_{s}\mathcal{H}[h_{s},\cdot,\cdot]=-\frac{\mu}{\rho^{2}}D^{3}\phi(z_{\rho})[\hat{h}_{s},\cdot,\cdot],$ (56b)
$D_{\hat{x}}\mathcal{H}^{-1}[h_{\hat{x}},\cdot,\cdot]=-\frac{\mu}{\rho}\mathcal{H}^{-1}D^{3}\phi(z_{\rho})[\hat{h}_{\hat{x}},\cdot,\cdot]\mathcal{H}^{-1},$ (56c)
$D_{s}\mathcal{H}^{-1}[h_{s},\cdot,\cdot]=\frac{\mu}{\rho^{2}}\mathcal{H}^{-1}D^{3}\phi(z_{\rho})[\hat{h}_{s},\cdot,\cdot]\mathcal{H}^{-1}.$ (56d)
@@
$D^{3}\eta_{\rho}(\hat{x},s;\mu)[h,h,h]$ (57)
$=$ $\ -\rho D_{\hat{x}}\mathcal{H}^{-1}[h_{\hat{x}},\mathcal{B}h_{\hat{x}},\mathcal{B}h_{\hat{x}}]-\rho D_{s}\mathcal{H}^{-1}[h_{s},h_{\hat{x}},h_{\hat{x}}$]
@@
$=$ $\ \mu D^{3}\phi(z_{\rho})[\hat{h}_{\hat{x}}-\rho^{-1}\hat{h}_{s},{\hat{h}_{\hat{x}}}-\rho^{-1}\hat{h}_{s},\hat{h}_{\hat{x}}-\rho^{-1}\hat{h}_{s}]$
$\overset{(i)}{\leq}$ ${2\mu}\left\{D^{2}\phi(z_{\rho})[\hat{h}_{\hat{x}}-\rho^{-1}\hat{h}_{s},\hat{h}_{\hat{x}}-\rho^{-1}\hat{h}_{s}]\right\}^{3/2}$
$\overset{(ii)}{\leq}$ $\ \dfrac{2}{\sqrt{\mu}}\left(S_{\eta_{\rho}}(\hat{x},s;\mu)[h,h]\right)^{3/2},$
where $(i)$ follows from the self-concordant property of $\phi$, and $(ii)$ follows from (54). Hence $\eta_{\rho}(\cdot,\cdot;\mu)$ is a nondegenerate $\mu$-self-concordant convex-concave function on $\hat{\mathbb{E}}\times\mathbb{E}$. The remaining statements follow immediately from Proposition 1.
3.3 Equivalent characterization of the parameterized smooth equations
This subsection is central to our analysis. We establish the equivalence between the parameterized smooth equations (25) and the first-order optimality conditions of the minimax problem (33).
@@ 3.3 Equivalent characterization of the parameterized smooth equations
For any $\mu>0$ and $\rho\geq 1$, the triple $(x(\mu),s(\mu),\lambda(\mu))$ solves the parameterized smooth equations (25) if and only if it is a primal-dual optimal solution of the barrier subproblem (58).
@@
By definition, the triple $(x(\mu),s(\mu),\lambda(\mu))$ solves the parameterized smooth equations (25) if and only if it satisfies
$\mathcal{A}x(\mu)-b=0,\;\mathcal{A}^{*}\lambda(\mu)+s(\mu)-c=0,\;\Phi_{\rho}(x(\mu),s(\mu);\mu)=0.$ (59)
@@
Consequently, $(x(\mu),s(\mu),\lambda(\mu))$ is a primal-dual optimal solution of the barrier subproblem (58). The converse implication follows by reversing the above arguments, and the proof is complete.
Based on Lemma 2, we now establish the equivalence between the parameterized smooth equations (25) and the first-order optimality conditions of the minimax problem (33).
Theorem 3.3
For any $\mu>0$ and $\rho\geq 1$, suppose that $(x(\mu),s(\mu),\lambda(\mu))$ solves the parameterized smooth equations (25). Then the pair $(\hat{x}(\mu),s(\mu)):=(\mathcal{B}^{*}(x(\mu)-\bar{x}),s(\mu))$ is a saddle point of the minimax problem (33). Conversely, if $(\hat{x}(\mu),s(\mu))$ is a saddle point of (33), then
solves the parameterized smooth equations (25).
Proof
Suppose that the triple $(x(\mu),s(\mu),\lambda(\mu))$ solves the parameterized smooth equations (25). Then by Lemma 2, we have
This combined with Lemma 1 yields
$z_{\rho}(x(\mu),s(\mu);\mu)=x(\mu).$ (62)
@@
$\nabla_{s}\eta_{\rho}(\hat{x}(\mu),s(\mu);\mu)=z_{\rho}(x(\mu),s(\mu);\mu)-x(\mu)=0.$ (63)
Moreover,
$\nabla_{\hat{x}}\eta_{\rho}(\hat{x},s;\mu)$ $=\mathcal{B}^{*}(c-y_{\rho}(x(\mu),s(\mu);\mu))$ (64)
$=\mathcal{B}^{*}(c-s(\mu)-\rho(z_{\rho}(x(\mu),s(\mu);\mu)-x(\mu)))$
$\overset{(i)}{=}\mathcal{B}^{*}(c-s(\mu))$
$\overset{(ii)}{=}\mathcal{B}^{*}\mathcal{A}^{*}\lambda(\mu)$
$=0,$
where $(i)$ follows from $z_{\rho}(x(\mu),s(\mu);\mu)=x(\mu)$, and $(ii)$ follows from the second equation in (61). Therefore, $(\hat{x}(\mu),s(\mu)):=(\mathcal{B}^{*}(x(\mu)-\bar{x}),s(\mu))$ satisfies the first-order optimality conditions of the minimax problem (33). Since $\eta_{\rho}$ is convex-concave, this pair is a saddle point of (33).
@@
Since $\mathcal{A}$ is surjective, and $\mathcal{A}^{*}(\mathcal{A}\mathcal{A}^{*})^{-1}\mathcal{A}$ is the orthogonal projector onto $({\rm ker}\,\mathcal{A})^{\perp}$, we obtain
$\mathcal{A}^{*}\lambda(\mu)+s(\mu)-c=(\mathcal{I}_{\mathbb{E}}-\mathcal{A}^{*}(\mathcal{A}\mathcal{A}^{*})^{-1}\mathcal{A})(s(\mu)-c)=0,$ (67)
@@
According to (Nesterov and Todd, 1998, Theorem 4.1), the barrier subproblem (58) admits a unique primal-dual optimal solution. This combined with Theorem 3.3 implies that the minimax problem (33) admits a unique saddle point for any $\mu>0$ and $\rho\geq 1$.
@@
Let $(x,s,\lambda)$ satisfy the linear constraint $\mathcal{A}x=b$. For any scalars $\mu>0$ and $\rho\geq 1$, suppose that $(\Delta x,\Delta s,\Delta\lambda)$ is the search direction generated by the SNM, satisfying (45). Then the pair $(\Delta\hat{x},\Delta s):=(\mathcal{B}^{*}\Delta x,\Delta s)$ is the unique Newton direction for the minimax problem (33), given by
@@
Suppose that $(\Delta x,\Delta s,\Delta\lambda)$ is the search direction generated by the SNM. Let $(\Delta\hat{x},\Delta s)=(\mathcal{B}^{*}\Delta x,\Delta s)$. Then,
$\rho\mathcal{B}^{*}\mathcal{H}^{-1}\mathcal{W}\mathcal{B}\Delta\hat{x}-\mathcal{B}^{*}\mathcal{H}^{-1}\mathcal{W}\Delta s+\mathcal{B}^{*}(c-y_{\rho})$ (70)
$\overset{(i)}{=}$ $\ \mathcal{B}^{*}(\rho\mathcal{H}^{-1}\mathcal{W}\Delta x-\mathcal{H}^{-1}\mathcal{W}\Delta s+c-s-\rho(z_{\rho}-x))$
$\overset{(ii)}{=}$ $\ \mathcal{B}^{*}(c-s-\Delta s)$
$\overset{(iii)}{=}$ $\ \mathcal{B}^{*}\mathcal{A}^{*}(\lambda+\Delta\lambda)$
$=$ $0,$
where $(i)$ follows from the definition of $y_{\rho}(x,s;\mu)$, $(ii)$ and $(iii)$ are the third and second equations of (45), respectively. The second equation of (68) follows directly from the third equation of (45). Consequently, the equations (68) admit at least one solution. Since the Schur complement of the operator in (68) relative to $\rho^{-1}\mathcal{H}^{-1}$ is $\rho\mathcal{B}^{*}\mathcal{W}\mathcal{B}\succ 0$, the uniqueness follows directly.
@@
$\mathcal{A}\Delta x=\mathcal{A}\mathcal{B}\Delta\hat{x}=0=-(\mathcal{A}x-b),$ (71)
which satisfies the first equation of (45). It suffices to verify the second equation. Left-multiplying the second equation in (68) by $\rho\mathcal{B}^{*}$ and adding it to the first equation, we obtain
$\mathcal{B}^{*}(c-s-\Delta s)=0.$ (72)
@@
$\mathcal{A}^{*}(\lambda+\Delta\lambda)+s+\Delta s-c$ $=(\mathcal{I}_{\mathbb{E}}-\mathcal{A}^{*}(\mathcal{A}\mathcal{A}^{*})^{-1}\mathcal{A})(s+\Delta s-c).$ (73)
@@
$\mathcal{A}^{*}(\lambda+\Delta\lambda)+s+\Delta s-c=0,$
which is the second equation of (45). The uniqueness of $(\Delta x,\Delta s,\Delta\lambda)$ follows from Proposition 2, and the proof is complete.
Theorems 3.3 and 3.4 provide an equivalent characterization of both the parameterized smooth equations and the search direction generated by the SNM under consideration via the reduced SBAL function $\eta_{\rho}$. This equivalence implies that, in the subsequent algorithmic analysis, it suffices to study the Newton iterations applied to the minimax problem (33). This offers a convenient and powerful tool for analyzing the behavior of the SNM. We conclude this section with a summary of the main characterizations obtained.

### 21 May 2026 v2

@@ 3.2 Properties of the reduced BAL function
$S_{\eta_{\rho}}(w;\mu)[h,h]=\rho\langle h_{\hat{x}},\mathcal{B}^{*}\mathcal{H}^{-1}\mathcal{W}\mathcal{B}h_{\hat{x}}\rangle+\rho^{-1}\langle h_{s},\mathcal{H}^{-1}h_{s}\rangle.$
Since $\mathcal{B}$ is injective and $\mathcal{H}^{-1}\mathcal{W}\succ 0$, the operator $\mathcal{B}^{*}\mathcal{H}^{-1}\mathcal{W}\mathcal{B}$ is positive definite on $\hat{\mathbb{E}}$. Combined with $\mathcal{H}^{-1}\succ 0$, this shows that $S_{\eta_{\rho}}(w;\mu)$ is positive definite, so $\eta_{\rho}(w;\mu)$ is nondegenerate. By Definition 3, it suffices to prove that
$|D^{3}\eta_{\rho}(w;\mu)[h,h,h]|\leq\dfrac{2}{\sqrt{\mu}}(S_{\eta_{\rho}}(w;\mu)[h,h])^{3/2}.$
Let $\bar{h}=(\bar{h}_{x},\bar{h}_{s})=(\mathcal{H}^{-1}\mathcal{B}h_{\hat{x}},\mathcal{H}^{-1}h_{s})\in\mathbb{E}\times\mathbb{E}$. By Corollary 1 and the formula $\mathcal{H}=\mathcal{I}_{\mathbb{E}}+\mathcal{W}=\mathcal{I}_{\mathbb{E}}+\frac{\mu}{\rho}D^{2}\phi(z_{\rho})$, we have
$S_{\eta_{\rho}}(w;\mu)[h,h]$ (45)
$=$ $\ \rho\langle\bar{h}_{x},\mathcal{H}\mathcal{W}\bar{h}_{x}\rangle+\rho^{-1}\langle\bar{h}_{s},\mathcal{H}\bar{h}_{s}\rangle$
$=$ $\ \mu D^{2}\phi(z_{\rho})[\bar{h}_{x},\bar{h}_{x}]+\frac{\mu^{2}}{\rho}\|D^{2}\phi(z_{\rho})\bar{h}_{x}\|^{2}+\rho^{-1}\|\bar{h}_{s}\|^{2}+\frac{\mu}{\rho^{2}}D^{2}\phi(z_{\rho})[\bar{h}_{s},\bar{h}_{s}]$
$=$ $\ \mu D^{2}\phi(z_{\rho})[\bar{h}_{x}-\rho^{-1}\bar{h}_{s},\bar{h}_{x}-\rho^{-1}\bar{h}_{s}]+\dfrac{1}{\rho}\|\mu D^{2}\phi(z_{\rho})\bar{h}_{x}+\bar{h}_{s}\|^{2}$
$\geq$ $\ \mu D^{2}\phi(z_{\rho})[\bar{h}_{x}-\rho^{-1}\bar{h}_{s},\bar{h}_{x}-\rho^{-1}\bar{h}_{s}],$
@@
$D^{2}\eta_{\rho}(w;\mu)[h,h]$ $=\rho\langle\mathcal{B}h_{\hat{x}},(\mathcal{I}_{\mathbb{E}}-\mathcal{H}^{-1})\mathcal{B}h_{\hat{x}}\rangle-\rho^{-1}\langle h_{s},\mathcal{H}^{-1}h_{s}\rangle$
@@
A direct calculation gives
$D_{\hat{x}}\mathcal{H}^{-1}[h_{\hat{x}},\cdot,\cdot]=-\frac{\mu}{\rho}D^{3}\phi(z_{\rho})[\bar{h}_{x},\mathcal{H}^{-1}\cdot,\mathcal{H}^{-1}\cdot],$ (46a)
$D_{s}\mathcal{H}^{-1}[h_{s},\cdot,\cdot]=\frac{\mu}{\rho^{2}}D^{3}\phi(z_{\rho})[\bar{h}_{s},\mathcal{H}^{-1}\cdot,\mathcal{H}^{-1}\cdot].$ (46b)
@@
$D^{3}\eta_{\rho}(w;\mu)[h,h,h]$
$=$ $\ -\rho D_{\hat{x}}\mathcal{H}^{-1}[h_{\hat{x}},\mathcal{B}h_{\hat{x}},\mathcal{B}h_{\hat{x}}]-\rho D_{s}\mathcal{H}^{-1}[h_{s},\mathcal{B}h_{\hat{x}},\mathcal{B}h_{\hat{x}}$]
@@
$=$ $\ \mu D^{3}\phi(z_{\rho})[\bar{h}_{x}-\rho^{-1}\bar{h}_{s},{\bar{h}_{x}}-\rho^{-1}\bar{h}_{s},\bar{h}_{x}-\rho^{-1}\bar{h}_{s}].$
Taking absolute values and using the standard self-concordance of $\phi$, we obtain
$\left|D^{3}\eta_{\rho}(w;\mu)[h,h,h]\right|=$ $\ \mu\left|D^{3}\phi(z_{\rho})[\bar{h}_{x}-\rho^{-1}\bar{h}_{s},{\bar{h}_{x}}-\rho^{-1}\bar{h}_{s},\bar{h}_{x}-\rho^{-1}\bar{h}_{s}]\right|$
$\leq$ ${2\mu}\left\{D^{2}\phi(z_{\rho})[\bar{h}_{x}-\rho^{-1}\bar{h}_{s},\bar{h}_{x}-\rho^{-1}\bar{h}_{s}]\right\}^{3/2}$
$\leq$ $\ \dfrac{2}{\sqrt{\mu}}\left(S_{\eta_{\rho}}(w;\mu)[h,h]\right)^{3/2},$
where the second inequality follows from (45). Hence, $\eta_{\rho}(\cdot,\cdot;\mu)$ is a nondegenerate $\mu$-self-concordant convex-concave function on $\hat{\mathbb{E}}\times\mathbb{E}$. The remaining statements follow immediately from Proposition 1.
3.3 Equivalent characterization of the parameterized smooth system
This subsection establishes the key equivalence between the parameterized smooth system (23) and the first-order optimality conditions of the minimax problem (29).
@@ 3.3 Equivalent characterization of the parameterized smooth system
For any $\mu>0$ and $\rho>0$, the triple $(x(\mu),s(\mu),\lambda(\mu))$ solves the parameterized smooth system (23) if and only if it is the KKT triple of the barrier subproblem (47).
@@
By definition, the triple $(x(\mu),s(\mu),\lambda(\mu))$ solves (23) if and only if it satisfies
$\mathcal{A}x(\mu)-b=0,\;\mathcal{A}^{*}\lambda(\mu)+s(\mu)-c=0,\;\Phi_{\rho}(x(\mu),s(\mu);\mu)=0.$
@@
$\mathcal{A}x(\mu)=b,\,\mathcal{A}^{*}\lambda(\mu)+s(\mu)-c=0,\,x(\mu),\,s(\mu)\in\operatorname{int}(\mathbb{K}),\,x(\mu)\circ s(\mu)$ $=\mu e.$
Consequently, $(x(\mu),s(\mu),\lambda(\mu))$ is the KKT triple associated with the barrier subproblem (47). The converse implication follows by reversing the above arguments, and the proof is complete.
Based on Lemma 2, we now establish the equivalence between the parameterized smooth system (23) and the first-order optimality conditions of the minimax problem (29).
Theorem 3.3
For any $\mu>0$ and $\rho>0$, suppose that $(x(\mu),s(\mu),\lambda(\mu))$ solves the parameterized smooth system (23). Then the pair $(\hat{x}(\mu),s(\mu)):=(\mathcal{B}^{*}(x(\mu)-\bar{x}),s(\mu))$ is a saddle point of the minimax problem (29). Conversely, if $(\hat{x}(\mu),s(\mu))$ is a saddle point of (29), then
solves the parameterized smooth system (23).
Proof
Suppose that the triple $(x(\mu),s(\mu),\lambda(\mu))$ solves (23). By Lemma 2, we have
@@
Together with Lemma 1, this yields
$z_{\rho}(x(\mu),s(\mu);\mu)=x(\mu).$
@@
$\nabla_{s}\eta_{\rho}(\hat{x}(\mu),s(\mu);\mu)=z_{\rho}(x(\mu),s(\mu);\mu)-x(\mu)=0.$
Furthermore,
$\nabla_{\hat{x}}\eta_{\rho}(\hat{x}(\mu),s(\mu);\mu)$ $=\mathcal{B}^{*}(c-y_{\rho}(x(\mu),s(\mu);\mu))$
$=\mathcal{B}^{*}(c-s(\mu)-\rho(z_{\rho}(x(\mu),s(\mu);\mu)-x(\mu))).$
By the identity $z_{\rho}(x(\mu),s(\mu);\mu)=x(\mu)$ and the second equation in (48), we have
$\nabla_{\hat{x}}\eta_{\rho}(\hat{x}(\mu),s(\mu);\mu)=\mathcal{B}^{*}(c-s(\mu))=\mathcal{B}^{*}\mathcal{A}^{*}\lambda(\mu)=0.$
Therefore, $(\hat{x}(\mu),s(\mu)):=(\mathcal{B}^{*}(x(\mu)-\bar{x}),s(\mu))$ satisfies the first-order optimality conditions of the minimax problem (29). Since $\eta_{\rho}$ is convex-concave, this pair is a saddle point of (29).
@@
Since $\mathcal{A}$ is surjective and $\mathcal{A}^{*}(\mathcal{A}\mathcal{A}^{*})^{-1}\mathcal{A}$ is the orthogonal projector onto $({\rm ker}\,\mathcal{A})^{\perp}$, we obtain
$\mathcal{A}^{*}\lambda(\mu)+s(\mu)-c=(\mathcal{I}_{\mathbb{E}}-\mathcal{A}^{*}(\mathcal{A}\mathcal{A}^{*})^{-1}\mathcal{A})(s(\mu)-c)=0,$
@@
According to (Nesterov and Todd, 1998, Theorem 4.1), the barrier subproblem (47) admits a unique primal-dual optimal solution. Together with Theorem 3.3, this implies that the minimax problem (29) admits a unique saddle point for any $\mu>0$ and $\rho>0$.
@@
Let $(x,s,\lambda)$ satisfy the affine constraint $\mathcal{A}x=b$. For any scalars $\mu>0$ and $\rho>0$, suppose that $(\Delta x,\Delta s,\Delta\lambda)$ is the search direction generated by the SNM, satisfying (40). Then the pair $(\Delta\hat{x},\Delta s):=(\mathcal{B}^{*}\Delta x,\Delta s)$ is the unique Newton direction for the minimax problem (29), given by
@@
Suppose that $(\Delta x,\Delta s,\Delta\lambda)$ is the search direction generated by the SNM. Let $(\Delta\hat{x},\Delta s)=(\mathcal{B}^{*}\Delta x,\Delta s)$. We first show that $\Delta x=\mathcal{B}\Delta\hat{x}$. Since $\mathcal{A}\Delta x=-(\mathcal{A}x-b)=0$, there exists a unique vector $\Delta\bar{x}\in\hat{\mathbb{E}}$ such that $\Delta x=\mathcal{B}\Delta\bar{x}$. It follows that
$\Delta\hat{x}=\mathcal{B}^{*}\Delta x=\mathcal{B}^{*}\mathcal{B}\Delta\bar{x}=\Delta\bar{x}.$
Thus, $\Delta\hat{x}=\Delta\bar{x}$ and $\Delta x=\mathcal{B}\Delta\hat{x}$. Consequently,
$\rho\mathcal{B}^{*}\mathcal{H}^{-1}\mathcal{W}\mathcal{B}\Delta\hat{x}-\mathcal{B}^{*}\mathcal{H}^{-1}\mathcal{W}\Delta s+\mathcal{B}^{*}(c-y_{\rho})$
$=$ $\ \mathcal{B}^{*}(\rho\mathcal{H}^{-1}\mathcal{W}\Delta x-\mathcal{H}^{-1}\mathcal{W}\Delta s+c-s-\rho(z_{\rho}-x))$
$=$ $\ \mathcal{B}^{*}(c-s-\Delta s)$
$=$ $\ \mathcal{B}^{*}\mathcal{A}^{*}(\lambda+\Delta\lambda)$
$=$ $0.$
Here the second and third equalities follow from the third and second equations of (40), respectively. The second equation of (50) follows directly from the third equation of (40). Consequently, the linear system (50) admits at least one solution. Since the Schur complement of the operator in (50) relative to $\rho^{-1}\mathcal{H}^{-1}$ is $\rho\mathcal{B}^{*}\mathcal{W}\mathcal{B}\succ 0$, the uniqueness follows directly.
@@
$\mathcal{A}\Delta x=\mathcal{A}\mathcal{B}\Delta\hat{x}=0=-(\mathcal{A}x-b),$
so the first equation of (40) holds. Furthermore, the second block row of (50) gives
$\mathcal{H}^{-1}\mathcal{W}\mathcal{B}\Delta\hat{x}-\rho^{-1}\mathcal{H}^{-1}\Delta s=-(z_{\rho}-x).$
Since $\Delta x=\mathcal{B}\Delta\hat{x}$, this is equivalent to
$\mathcal{H}^{-1}\mathcal{W}\Delta x+\rho^{-1}\mathcal{H}^{-1}\Delta s=z_{\rho}-x,$
which is precisely the third equation of (40). It remains to verify the second equation of (40). Multiplying the second equation in (50) on the left by $\rho\mathcal{B}^{*}$ and adding it to the first equation, we obtain
$\mathcal{B}^{*}(c-s-\Delta s)=0.$
@@
$\mathcal{A}^{*}(\lambda+\Delta\lambda)+s+\Delta s-c$ $=(\mathcal{I}_{\mathbb{E}}-\mathcal{A}^{*}(\mathcal{A}\mathcal{A}^{*})^{-1}\mathcal{A})(s+\Delta s-c).$
@@
$\mathcal{A}^{*}(\lambda+\Delta\lambda)+s+\Delta s-c=0.$
Thus all three equations in (40) are satisfied. The uniqueness of $(\Delta x,\Delta s,\Delta\lambda)$ follows from Proposition 2, and the proof is complete.
Theorems 3.3 and 3.4 provide an equivalent characterization of both the parameterized smooth system and the search direction generated by the SNM via the reduced BAL function $\eta_{\rho}$. This equivalence implies that, in the subsequent algorithmic analysis, it suffices to study the Newton iterations applied to the minimax problem (29). This offers a convenient and powerful tool for analyzing the behavior of the SNM. We conclude this section with a summary of the main characterizations obtained.